OpenStudy (clairexiaoke):

If the instructions for a problem ask you to use the smallest possible domain to completely graph two periods of y = 5 + 3 cos 2(x - pi divided by three), what should be used for Xmin and Xmax? Explain your answer.

OpenStudy (clairexiaoke):

@satellite73

OpenStudy (clairexiaoke):

the equation is y=5+3 cos2 (x-(pi/3))

OpenStudy (clairexiaoke):

@.Sam.

OpenStudy (clairexiaoke):

@campbell_st

OpenStudy (thomas5267):

Is \(y=5+3\cos\left(2x-\frac{2\pi}{3}\right)\)?

OpenStudy (clairexiaoke):

\[y=5+3\cos 2(x-\frac{ \pi }{ 3 })\]

OpenStudy (anonymous):

thats better

OpenStudy (anonymous):

of course it is the same thing that @thomas5267 wrote

OpenStudy (clairexiaoke):

i suppose you'll have to find the domain first right?

OpenStudy (anonymous):

this is two questions right ?

OpenStudy (clairexiaoke):

it's asking for Xmax and Xmin and you'll need to explain

OpenStudy (anonymous):

one is two periods where you have pretty much an infinite number of choices, since that is what preiodic means

OpenStudy (anonymous):

Xmax is not a math term i am familiar with

OpenStudy (anonymous):

ooh i see !!

OpenStudy (clairexiaoke):

i think it means x maxima

OpenStudy (anonymous):

it wants the left and right hand endpoints is all

OpenStudy (clairexiaoke):

and how do you solve for that?

OpenStudy (anonymous):

the easiest way for me would be to do this the period of cosine is \(2\pi\) so two periods would have length \(4\pi\)

OpenStudy (clairexiaoke):

yup i get that part

OpenStudy (clairexiaoke):

but how do i find the smallest possible domain?

OpenStudy (anonymous):

set \[2x-\frac{2\pi}{3}=0\] solve for \(x\) then set \[2x-\frac{2\pi}{3}=4\pi\] and solve that for x

OpenStudy (clairexiaoke):

ok hold up

OpenStudy (clairexiaoke):

would x be \[\frac{ 2\pi }{ 6 }\] ?

OpenStudy (anonymous):

i would write \(\frac{\pi}{3}\) but yes

OpenStudy (clairexiaoke):

so \[4\pi\] is equal to 0

OpenStudy (anonymous):

what does \(4\pi=0\) mean?

OpenStudy (clairexiaoke):

oops. nvm i misread something

OpenStudy (clairexiaoke):

and the other one would be \[\frac{ 7p }{ 3 }\]

OpenStudy (clairexiaoke):

would that be the start point and the end point?

OpenStudy (anonymous):

that is one way

OpenStudy (anonymous):

but look, cosine is periodic

OpenStudy (anonymous):

you want two periods you can start and end anywhere \(4\pi\) units apart

OpenStudy (thomas5267):

I would add a random multiples of pi to both the start point and end point just to annoy the teacher lol.

OpenStudy (clairexiaoke):

but they're asking for the smallest domain which is why we started at 0 right?

OpenStudy (anonymous):

i guess i cannot get my point across you can start at 0 and went to \(4\pi\) because it is easy but you could go from \(-2\pi\) to \(2\pi\) instead

OpenStudy (anonymous):

for that matter you could go from \(-e^2\) to \(-e^2+4\pi\) it really makes no difference

OpenStudy (clairexiaoke):

yea i get that

OpenStudy (clairexiaoke):

then what do they mean by using the smallest domain, because that the part where it keeps throwing me off

OpenStudy (anonymous):

that is what we did

OpenStudy (anonymous):

they said interval of 4 pi and we did it

OpenStudy (thomas5267):

I guess smallest means xmax-xmin is the smallest?

OpenStudy (clairexiaoke):

oh ok, i think i get it now. thank you guys! :)

OpenStudy (clairexiaoke):

this helped a lot

OpenStudy (anonymous):

yw